News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Wednesday 10 May 2000 12.00pm EDT

During last week the observing schedule was replanned to include two observations of the target XTE J1550-564 which was accepted as a Target Of Opportunity (TOO) on 4/30 following a major X-ray flux increase detected by the ASM of RXTE. The revised schedule from May 3 on:

---------------------------------------------
XTE J1550-564 TOO      ACIS-S/HETG     May 3
Radiation Belts        None
CL 0848.6+4453         ACIS-I          May 4
NGC 4261               ACIS-S          May 5
XTE J1550-564 TOO      ACIS-S/HETG     May 6
Radiation Belts        None
SN 1993J               ACIS-S          May 7
M82                    ACIS-I
Proxima Centauri       ACIS-S/HETG
Abell 1995             ACIS-S          May 8
---------------------------------------------

An anomaly related to the High Energy Transmission Grating (HETG) was detected on 5/8 after the grating did not insert when commanded by the daily load. Telemetry from the previous retraction on 5/6 showed that the grating had retracted but had not triped a limit switch near the end of travel. The limit switch changes the state of a bi-level flag that is checked during the software start-up prior to any insertion. If the flag is not set to "retract" the software will not insert the HETG. The data and event time-line point to a faulty A-side limit switch as the cause.

Real-time procedures were run at 1pm EDT 5/8 and 6.30pm EDT 5/9 to collect additional A-side data and to verify that the B-side electronics and limit switch function as expected. The daily load for the remainder of the week was replanned to include imaging observations only, and uplinked to the spacecraft at 6.30pm EDT 5/9. An additional test is planned to complete B-side checkout before resuming HETG observations from the daily load. Work has begun to develop an approach to restore A-side redundancy, e.g., by using potentiometer data in place of the limit-switch state to signal the HETG retraction state.

Real-time procedures were also run on 5/1/00 to perform the third ACIS squeegee mode test, and on 5/4/00 to perform a routine Ephemeris Coefficient Update.

All spacecraft subsystems, with the exception of HETG as above, continued to operate nominally.





Return to Status Reports